OpenStudy (anonymous):

Simplify SQRT704 ? Not UnderStanding , Please Help .

OpenStudy (amistre64):

break down 704 into its prime factors

OpenStudy (amistre64):

or at least start hacking away at it by seeing what perfect squares will factor into it

OpenStudy (amistre64):

4,9,16,25,36, ... are perfect squares

OpenStudy (amistre64):

704/2 = 352 352/2 = 176 176/2 = 88 88/2 = 44 44/2 = 22 22/2 = 11 2.2.2.2.2.2.11 = 704 right?

OpenStudy (anonymous):

\[704=2^6\times 11\] because \[704=2\times 352=2^2\times 176=2^3\times 88=2^4\times 44=2^5\times 22\] \[=2^6\times 11\]

OpenStudy (anonymous):

But Why Are You Dividing By 2 ?

OpenStudy (anonymous):

good morning!

OpenStudy (amistre64):

becasue 2 is an easy number to divide by; it just whittles away at it till it gets small enough to see

OpenStudy (anonymous):

also because you CAN divide by 2. if you have 75 you would divide by 3 and 5

OpenStudy (anonymous):

divide by whatever number goes in to it until you have factored into the product of primes.

OpenStudy (amistre64):

right, if you cant divide by 2, work your way up to 3 and 5 and 7 etc

OpenStudy (amistre64):

now that its all factored, lets start pairing them up to move out of the radical. since we have a square root; we want to group in 2s, like noahs ark sqrt[(2.2).(2.2).(2.2).11] ; re work the math sqrt[(4).(4).(4).11] ; re work again sqrt(4).sqrt(4).sqrt(4).sqrt(11) ; re work again 2.2.2.sqrt(11) ; and one final try to get :) 6 sqrt(11)

OpenStudy (anonymous):

typo

OpenStudy (amistre64):

8 sqrt(11) :)

OpenStudy (amistre64):

dont forget to remember how to multiply lol

OpenStudy (anonymous):

\[\sqrt{704}=\sqrt{2^6\times 11}=2^3\times \sqrt{11}\]

OpenStudy (anonymous):

yah division is easier

OpenStudy (saifoo.khan):

Hi Satellite & A.Mirstre.

OpenStudy (anonymous):

saifooooooooooooooooooooooooooo!

OpenStudy (saifoo.khan):

Great Morning.

OpenStudy (saifoo.khan):

xD

OpenStudy (anonymous):

? afternoon right?

OpenStudy (saifoo.khan):

here its 16:54pm.